OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)? 1/5 -1/5 5 -5

OpenStudy (anonymous):

-1/5

OpenStudy (anonymous):

m = (y2-y1) / (x2-x1)

OpenStudy (johnweldon1993):

\[slope = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] You have 2 points (1, 5) and (6, 4)? (x1,y1) (x2,y2) so we can plug those numbers into that equation \[slope = \frac{ 4 - 5 }{ 6 - 1 }\] \[slope = -\frac{ 1 }{ 5 }\]